OpenStudy (anonymous):

Which polynomial identity will prove that 117 = 125 - 8?

OpenStudy (anonymous):

A Difference of Squares B Difference of Cubes C Sum of Cubes D Square of a Binomial

OpenStudy (anonymous):

Hint: \(125=5^3\) and \(8=2^3\).

OpenStudy (anonymous):

I was about to post that, but had a message to reply to:)

OpenStudy (anonymous):

how do you do these numbers inside?

OpenStudy (anonymous):

`\(125=5^3\) and \(8=2^3\)` Wrap your expressions in `\(` and `\)` for inline tex display.

OpenStudy (anonymous):

Difference of cubes?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

yes, Leah, siths, tnx:)

OpenStudy (anonymous):

Yay thank you guys!